# RouterBench: A Benchmark for Multi-LLM Routing System

### v1

@@ 5.4 RAG Results
Building on results above, we simultaneously compared various router types, including predictive and cascading routers, on the RAG dataset. We used the same setting for KNN routers and MLP routers while selecting error rate $0.2$ for cascading routers. We randomly partitioned RAG dataset into two splits: 70% for training predictive routers and 30% for evaluating all routers. Figure 6 demonstrates that all routers performed comparably to the Zero Router. Nevertheless, further analysis revealed that routers with internet access significantly outperformed GPT-4 and GPT-3.5 Turbo on news platform data due to their real-time information retrieval capability. Specifically, the KNN router demonstrated a preference for routing news data to models with internet access, although it struggled with wiki data, leading to sub-optimal overall results. Additional relevant visualizations are provided in Appendix D. Our findings highlight the potential of model routing to enhance LLM applications within the ”Compound AI Systems” [^4] https://bair.berkeley.edu/blog/2024/02/18/compound-ai-systems/ scenario.
Contrary to previous findings, the cascading router underperformed in this experiment, likely due to increased cost discrepancies among models and the inefficacy of starting with the least expensive model. This underscores the necessity for optimized LLM candidate ranking in cascading router designs, a principle that has also been demonstrated in (Chen et al., 2023).
Figure 6: Left: Total Cost vs Performance for five models and four routers on RAG dataset. The AIQ values are also calculated for all four routers. NDCH represents non-decreasing convex hull. Right: Individual model performances on data curated from Wikipedia and news, and KNN router selection frequencies on them respectively.
@@ 6 Limitations and Future Work
RouterBench currently only focuses on performance and economic cost. It is meaningful to include more evaluation criteria, such as latency, throughput, and others to capture a more comprehensive understanding of router capabilities and limitations. There are also many LLMs and tasks that are not included in RouterBench due to the limitation of time, and future iterations of this benchmark would include datasets that cover more tasks to effectively evaluate the ever-growing capability of LLMs, and also to add newer LLMs as they are being released.
Our current work only evaluates the efficacy of predictive and cascading routers, yet there exists considerable scope for investigating further router designs, as highlighted in Section 5.3. Delving into more advanced router designs is crucial for enhancing routing efficiency. It’s noteworthy that our evaluation within the RAG context was limited to models possessing inherent retrieval capabilities. Addressing the challenge of implementing two-stage routing, which encompasses both retrievers and LLMs, remains critical. Such an approach could significantly refine router evaluations on standard RAG tasks, including HotpotQA (Yang et al., 2018) and NaturalQuestions (Kwiatkowski et al., 2019), by ensuring more precise assessments.
Furthermore, although the seven datasets in RouterBenchoffer broad coverage across a variety of tasks, incorporating domain-specific tasks that require long-tail skills, like translation of low-resource languages, could reveal additional intriguing aspects of LLM routing. This enhancement would align the benchmark more closely with real-world application scenarios. Efforts to integrate such tasks in future versions are planned.
@@ 7 Conclusion
We present RouterBench, a benchmark specifically designed for the evaluation of router mechanisms within multi-LLM systems. By addressing the critical need for standardized evaluation in this domain, our benchmark provides a comprehensive dataset and a theoretical framework specifically designed for the nuanced analysis of router cost-efficiency and performance. The insights from our study shed light on the effectiveness of various routing strategies, emphasizing the necessity for advanced LLM routing systems and refined routing methods. This work establishes a robust and scalable benchmark for router evaluation and aims to facilitate future progress in the efficient and cost-effective deployment of Large Language Models.
@@ A.1 Model Details & Cost Estimation
A.2 Dataset Generation for RAG
We designed two subtasks of different nature in dataset generation step to capture the complexity and heterogeneity of real-work RAG use cases.
  • Wikipedia: this subtask consists of 308 questions generated from wikipedia pages that focus on historic perspective or timeless content (e.g. wikipedia pages on history of pizza, or black holes).
  • News: this subtask consists of 349 questions generated from news articles of major new outlets published within 1 month of experiment.
Different LLM and compound AI systems are expected to have different performance on those subtasks of different nature, and thus create room for routing.
A.3 Dataset Details
@@
Hellaswag (Zellers et al., 2019): A dataset that challenges models to pick the best ending choice to a sentence given. It uses Adversarial Filtering(AF) to create a Goldilocks zone of complexity, wherein generations are largely nonsensical to humans but always make models struggle.
@@
MBPP (Austin et al., 2021): The benchmark is designed to be solvable by entry level programmers, covering programming fundamentals, standard library functionality, and so on. Each problem consists of a task description, code solution and 3 automated test cases.
MT-Bench (Zheng et al., 2023b): This dataset contains 3.3K expert-level pairwise human preferences for model responses generated by 6 models in response to 80 MT-bench questions, multi-run QA. The 6 models are GPT-4, GPT-3.5888https://openai.com/blog/chatgpt, Claude-v1, Vicuna-13B (Zheng et al., 2023a), Alpaca-13B (Taori et al., 2023), and LLaMA-13B (Touvron et al., 2023). The annotators are mostly graduate students with expertise in the topic areas of each of the questions. In this work, we only used the 80 questions to generate model responses for RouterBench.
A.4 More Details on Dataset Construction
@@
  • $eval\_name$: the source data this specific sample comes from. Example: hellaswag.dev.v0
  • $prompt$: prompt sentence. Example: The following are multiple choice questions…
@@
  • $performance$: the result compared to true label. Example: True/False
@@
A.5 Evaluation Metrics
We will perform 5-shot inference on MMLU, HellaSwag, GSM8K, ARC Challenge, Winogrande and 0-shot inference on MBPP, MT-Bench and RAG.
@@ A.5 Evaluation Metrics
A.6 Individual Dataset Result
@@ Appendix C Issues with Overly-aligned Models
Appendix D More Results on RAG Data
In this section, additional visualizations of the RAG dataset outcomes are presented. Figure 8 highlights the variability in router performance across the Wikipedia and news segments of the RAG dataset, resulting in less-than-ideal overall outcomes. Concurrently, Figure 9 reveals that the embeddings of Wikipedia and news platform samples exhibit distinct characteristics, suggesting they are distinguishable. This observation underscores the significant potential for enhancing router designs to elevate performance on this task in our RouterBench.
Figure 8: Total Cost vs Performance for five models and four routers on the Wikipedia and news platforms data partitions in the RAG dataset.
Figure 9: t-SNE visualization for wiki and news data embeddings.
Appendix E Full Cascading Routers Results
@@
Appendix F Training Data Distribution

### v2

@@ 5.4 RAG Results
Building on the results above, we simultaneously compared various router types, including predictive and cascading routers, on the RAG dataset. We used the same setting for KNN and MLP routers while selecting an error rate 0.2 for cascading routers. We randomly partitioned the RAG dataset into two splits: 70% for training predictive routers and 30% for evaluating all routers. Figure 6 demonstrates that all routers significantly improve compared to the Zero Router. Further analysis shows that the routers can identify time-sensitive features (like ”2024”) in user queries and route to online models for time-sensitive queries and GPT-4/GPT-3.5 for time-insensitive queries. Our findings highlight the potential of model routing to enhance LLM applications within the ”Compound AI Systems” (Zaharia et al., 2024) scenario.
Figure 6: Total Cost vs Performance for five models and four routers on the RAG dataset. The AIQ values are also calculated for all four routers. NDCH represents the non-decreasing convex hull.
@@ 6 Limitations and Future Work
RouterBench currently only focuses on performance and economic cost. It is meaningful to include more evaluation criteria, such as latency, throughput, and others, to capture a more comprehensive understanding of router capabilities and limitations. There are also many LLMs and tasks that are not included in RouterBench due to the limitation of time, and future iterations of this benchmark would include datasets that cover more tasks to evaluate the ever-growing capability of LLMs effectively and also to add newer LLMs as they are being released.
Our current work only evaluates the efficacy of predictive and cascading routers, yet considerable scope exists for investigating further router designs, as highlighted in Section 5.3. Delving into more advanced router designs is crucial for enhancing routing efficiency. Notably, our evaluation within the RAG context was limited to models possessing inherent retrieval capabilities. Addressing the challenge of implementing two-stage routing, which encompasses retrievers and LLMs, remains critical. This approach could significantly refine router evaluations on standard RAG tasks, including HotpotQA (Yang et al., 2018) and NaturalQuestions (Kwiatkowski et al., 2019), by ensuring more precise assessments.
Furthermore, although the seven datasets in RouterBenchoffer broad coverage across various tasks, incorporating domain-specific tasks that require long-tail skills, like translation of low-resource languages, could reveal additional intriguing aspects of LLM routing. This enhancement would align the benchmark more closely with real-world application scenarios. Efforts to integrate such tasks in future versions are planned.
@@ 7 Conclusion
We present RouterBench, a benchmark specifically designed to evaluate routers for multi-LLM systems. By addressing the critical need for standardized evaluation in this domain, our benchmark provides a comprehensive dataset and a theoretical framework designed for the nuanced analysis of router cost-efficiency and performance. The insights from our study shed light on the effectiveness of various routing strategies and revealed promising early results in some tasks. This work establishes a robust and scalable benchmark for router evaluation and aims to facilitate future progress in the efficient and cost-effective deployment of Large Language Models.
@@ A.1 Model Details & Cost Estimation
A.2 Dataset Details
@@
Hellaswag (Zellers et al., 2019): This dataset challenges models to pick the best ending choice for a given sentence. It uses Adversarial Filtering(AF) to create a Goldilocks zone of complexity, wherein generations are largely nonsensical to humans but always make models struggle.
@@
MBPP (Austin et al., 2021): The benchmark is designed to be solvable by entry-level programmers, covering programming fundamentals, standard library functionality, etc. Each problem comprises a task description, code solution, and 3 automated test cases.
MT-Bench (Zheng et al., 2023b): This dataset contains 3.3K expert-level pairwise human preferences for model responses generated by 6 models in response to 80 MT-bench questions, multi-run QA. The 6 models are GPT-4, GPT-3.5777https://openai.com/blog/chatgpt, Claude-v1, Vicuna-13B (Zheng et al., 2023a), Alpaca-13B (Taori et al., 2023), and LLaMA-13B (Touvron et al., 2023). The annotators are mostly graduate students with expertise in the topic areas of each of the questions. In this work, we only used the 80 questions to generate model responses for RouterBench.
A.3 More Details on Dataset Construction
@@
  • $eval\_name$: the source data from which this specific sample comes. Example: hellaswag.dev.v0
  • $prompt$: prompt sentence. Example: The following are multiple-choice questions…
@@
  • $performance$: the result compared to the true label. Example: True/False
@@
A.4 Evaluation Metrics
We will perform 5-shot inference on MMLU, HellaSwag, GSM8K, ARC Challenge, Winogrande, and 0-shot inference on MBPP, MT-Bench, and RAG.
@@ A.4 Evaluation Metrics
A.5 Individual Dataset Result
@@ Appendix C Issues with Overly-aligned Models
Appendix D Full Cascading Routers Results
@@
Appendix E Training Data Distribution